OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. x^2 - 4x + y^2 + 8y = -4 Someone please help me?!

OpenStudy (anonymous):

@Mertsj Could you help? i would appreciate it

OpenStudy (anonymous):

@ranga @Goldenmoon @Gabylovesyou @Philippians413 @agent0smith @hba

OpenStudy (anonymous):

@helpme1.2 @Captainskrullz @vorobya @britneyspears @nikato @mahone331 Help me out?

OpenStudy (nikato):

Do u know how to complete the square?

OpenStudy (anonymous):

No i dont understand this at all.

OpenStudy (nikato):

The equation of a circle is usually written in the form (x-h)^2 + (y-k)^2 = r^2 Where h,k is the center and r is the radius

OpenStudy (anonymous):

ohh okay

OpenStudy (nikato):

So ur problem is (x^2-4x+_) +(y^2+8x+_)=-4

OpenStudy (nikato):

Ur goal now is to try to fill in those 2 blanks

OpenStudy (anonymous):

okay.

OpenStudy (nikato):

And to fill in the blank U have to complete the square Lets start with x^2-4x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

how would i do that?

OpenStudy (nikato):

So u know how quadratic eqauionnr written in ax^2+bx+c

OpenStudy (anonymous):

yeah okay.

OpenStudy (nikato):

When u complete the square, u want to find c. To find c, u divide b by 2 and then square it

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@nikato could you please continue?

OpenStudy (nikato):

Ok, so the c in x^2-4x is 4. Remember the c is always positive. And u'll get (x^2-4x+4) This is how I got the 4. The b was -4. So -4/2=-2 and (-2)^2 =4 So can u try to find the c for y^2+8y?

OpenStudy (anonymous):

wait so the blank in the equation x^2 - 4x is 4?

OpenStudy (nikato):

+4 And now can u try to find the second blank?

OpenStudy (anonymous):

is it 8?

OpenStudy (nikato):

No, so what's the b?

OpenStudy (anonymous):

b is -4 right?

OpenStudy (nikato):

No, it's 8. y^2+8x and b would be 8 To find c, divide 8 by 2 and then square ur answer

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

Now what? what is the center and radius?

OpenStudy (nikato):

Wait, so what did u get for c

OpenStudy (anonymous):

i got 8

OpenStudy (nikato):

No, it's 16 8/2=4 and 4^2=16

OpenStudy (anonymous):

ohh yeah square! Oops

OpenStudy (nikato):

So now u have (x^2-4x+4)+(y^2+8x+16)=-4+4+16 Added 4 and 16 which are ur c. Becuz u added them on ur left side u have to add them to the right also

OpenStudy (nikato):

Factor (x-2)^2 +(y+5)^2 = 4^2

OpenStudy (nikato):

And if u look back at one of my earlier post, u can see which is ur center and raius

OpenStudy (anonymous):

Is the center (2,5) and the radius 4?

OpenStudy (nikato):

Center is (2,-5) and ur radius is correct

OpenStudy (anonymous):

Okay thank you so much!

OpenStudy (nikato):

Yea, no problem